# Stochastic Thermodynamics of Non-reciprocally Interacting Particles and Fields

### v2

@@ Appendix A Microscopic excess and house-keeping decomposition of the EPR
The total mean microscopic bulk EPR can further be divided using another physical aspect of the system. For example, when a state variable is easier to track rather than the underlying thermodynamic energy functional. One particular case is: when the temporal dynamics of the microstate leads to a steady-state probability distribution $P_{\{N\}}^{ss}$ for the microstate $\{N\}$. The mean bulk EPR ($\langle\dot{\sigma}^{\mathcal{R}}\rangle+\langle\dot{\sigma}^{\mathcal{D}}\rangle$) can be divided into two components called excess and housekeeping EPR using the steady state of the dynamics [293, 233, 294, 20, 99, 100, 232]. This decomposition is usually referred to as the Hatano-Sasa (HS) decomposition of the total mean EPR. The mean excess HS EPR is denoted by $\langle\dot{\sigma}^{ex}_{hs}\rangle$. The exact expression for $\langle\dot{\sigma}^{ex}_{hs}\rangle$ reads:
@@
Using the master eq. 7, the mean microscopic excess HS EPR satisfies, $\langle\dot{\sigma}^{ex}_{hs}\rangle=\sum_{\{N\}}d_{t}P_{\{N\}}(t)\ln{\left(P^{ss}_{\{N\}}(t)\right)}$. The time dependence of the steady state denotes the validity of the definition eq. 71 for the non-autonomous dynamics implemented through the thermodynamic work. The exact expression for the microscopic housekeeping EPR reads:
@@
The positivity of excess and housekeeping EPR has been shown [233, 294, 99, 100, 232]. Combining the rate of change of Gibbs entropy term $d_{t}s^{gb}$ with the mean excess HS EPR $\langle\dot{\sigma}^{ex}_{hs}\rangle$, one defines the non-adiabatic mean EPR $\langle\dot{\sigma}^{na}\rangle$. The closed-form expression for $\langle\dot{\sigma}^{na}\rangle=\langle\dot{\sigma}^{ex}_{hs}\rangle+d_{t}s^{gb}$ reads:
@@
The non-adiabatic mean EPR $\langle\dot{\sigma}^{ex}_{hs}\rangle$ quantifies the microscopic EPR due to a mismatch between the instantaneous probability distribution $\{P_{\{N\}}(t)\}$ and the steady state probability distribution $\{P^{ss}_{\{N\}}(t)\}$. In steady state the instantaneous probability distribution $\{P_{\{N\}}(t)\}=\{P^{ss}_{\{N\}}(t)\}$ is satisfied. Thus in steady-state the non-adiabatic mean EPR $\langle\dot{\sigma}^{na}\rangle$ vanishes. In steady state, the total mean microscopic EPR $\langle\dot{\sigma}\rangle$ is completely determined by the mean housekeeping EPR $\langle\dot{\sigma}^{hk}_{hs}\rangle$, such that $\langle\dot{\sigma}^{hk}_{hs}\rangle=\langle\dot{\sigma}\rangle$ holds. Reorganising the definition for the mean microscopic excess HS EPR and the microscopic non-adiabatic EPR it can be observed that they are related to the rate of change of stochastic Massieu Potential. In particular, $\langle\dot{\sigma}^{ex}_{hs}\rangle=d_{t}\langle\ln{(P^{ss}_{\{N\}}(t))}\rangle$ and $\langle\dot{\sigma}^{na}\rangle=-d_{t}\langle\ln{(P_{\{N\}}(t))/(P^{ss}_{\{N\}}(t))}\rangle$. Where the time derivative is defined as $d_{t}\langle*\rangle=\sum_{\{N\}}(*)d_{t}P_{\{N\}}$, the same convention has previously been utilized in defining the second law of thermodynamics eq. 13. Thus $-\ln{(P^{ss}_{\{N\}}(t))}$ and $\ln{(P_{\{N\}}(t))/(P^{ss}_{\{N\}}(t))}$ are the stochastic Massieu potential whose temporal variation is $\langle\dot{\sigma}^{ex}_{hs}\rangle$ and $\langle\dot{\sigma}^{na}\rangle$ respectively.
@@
The Kullback-Leibler divergence is defined as $D^{KL}_{ss}(P(t)||P^{ss}(t))=\sum_{\{N\}}P_{\{N\}}(t)\ln{(P_{\{N\}}(t)/P^{ss}_{\{N\}}(t))}$ [94, 95, 97, 98, 96, 99, 100, 101]. We introduce a shorthand notation for the $D^{KL}_{ss}(t)=D^{KL}(P(t)||P^{ss}(t))$, superscript gives index for the reference probability distribution. Thus, $\langle\dot{\sigma}^{na}\rangle=-d_{t}D^{KL}_{ss}(P(t))$ is satisfied. Physically it relates the non-adiabatic EPR to the relaxation of the Kullback-Leibler divergence between the instantaneous and steady-state probability distributions. Plugging this into eq. 72 and using eq. 13, it can be deduced that $\langle\dot{\sigma}^{hk}_{hs}\rangle=\langle\dot{\sigma}\rangle|_{P^{ss}}$. Hence, housekeeping EPR is the total microscopic EPR in steady-state and the microscopic adiabatic EPR vanishes. For a system without an external non-conservative driving force or non-reciprocal force, $\langle\dot{\sigma}^{hk}_{hs}\rangle=0$ is trivially satisfied. The microscopic EPR due to the relaxation process is fully quantified by the rate of change of the Kullback-Leibler divergence, thus $\langle\dot{\sigma}\rangle=-d_{t}D^{KL}_{ss}(t)$. A stronger bound on the microscopic EP is obtained by identifying the positivity of the $\langle\dot{\sigma}^{na}\rangle$ and the Kullback-Leibler divergence representation of it [101]. In particular, $\langle\sigma^{na}\rangle=D^{KL}_{ss}(0)-D^{KL}_{ss}(\tau)$, here $\tau$ is the relaxation time. It can further be simplified due to the triangle inequality [101, 98], namely, $\langle\sigma^{na}\rangle\geq D^{KL}(P(0)||P(\tau))$, which is tighter than the second law of the thermodynamics due to the non-negativity of the KL divergence.
Appendix B Information thermodynamics
The framework of information thermodynamics has been developed that takes into account the thermodynamic implications of the statistical information. The statistical information about the system state is encoded in Gibbs’s entropy of the system. One can achieve an information gain through a feedback-controlled operation or measurement [295, 296, 231, 297, 298]. This information gain reduces the statistical uncertainty associated with the system resulting in a lower EP signifying the reduced statistical uncertainty.
We denote the hidden variable by $Y$ that does not affect the transition dynamics (bulk term) of the macrostate/mesostate. The stochastic mutual information $I_{\{\rho\},Y}$ is defined as:
We have used the Bayes theorem to reach the second equality in eq. 74. Mutual information quantifies the statistical correlation (or dependence) between $\{\rho\}$ and the hidden variable $Y$. Mutual information vanishes if the system macrostate/mesostate is statistically independent of the hidden variable, $\mathcal{P}_{\{\rho\}|Y}=\mathcal{P}_{\{\rho\}},\forall,Y\in\Omega_{Y},\{\rho\}\in\Omega_{\{\rho\}}$. Where $\Omega_{Y}$ and $\Omega_{\{\rho\}}$ denote the probability measure space for the variable $Y$ and macrostate / mesostate $\{\rho\}$, respectively. The mean mutual information is defined as, $\langle I_{\{\rho\},N}\rangle=\int_{\Omega_{Y}}\int_{\Omega_{\{\rho\}}}I_{\{\rho\},N}\mathcal{P}_{\{\rho\},Y}$. Where, $\int\mathbb{D}[\{\rho\}]\equiv\int_{\Omega_{\{\rho\}}}$ is the shorthand notation used for the probability measure space corresponding to the macrostate $\{\rho\}$. The mutual information EPR is defined as $\langle\dot{I}_{\{\rho\},N}\rangle=\int_{\Omega_{Y}}\int_{\Omega_{\{\rho\}}}I_{\{\rho\},N}d_{t}\mathcal{P}_{\{\rho\},Y}$. The exact expressions for the Mutual information EP and EPR reads:
$\begin{split}\langle I_{\{\rho\},Y}\rangle&=\mathcal{S}^{gb}(\mathcal{P}_{\{\rho\}|Y})-\mathcal{S}^{gb}(\mathcal{P}_{\{\rho\}}),\\ \langle\dot{I}_{\{\rho\},Y}\rangle&=d_{t}\mathcal{S}^{gb}(\mathcal{P}_{\{\rho\}|Y})-d_{t}\mathcal{S}^{gb}(\mathcal{P}_{\{\rho\}})\end{split}$ (75)
Incorporating the mutual information EP and EPR through, the second law of thermodynamics eq. 40 is modified. Such that, the Gibbs EPR is replaced by the conditional Gibbs EPR. The conditional Gibbs entropy satisfy the inequality $\mathcal{S}^{gb}(\mathcal{P}_{\{\rho\}})\geq\mathcal{S}^{gb}(\mathcal{P}_{\{\rho\}|Y})$. Physically, this corresponds to the information gained due to the hidden variable $Y$, which reduced the statistical ignorance (Gibbs entropy) of the state of the system by encapsulating its correlations with the hidden variable. $\langle\dot{I}_{\{\rho\},Y}\rangle\leq 0$ holds, which gives the sign convention for the information gain. The eq. 75 satisfies the following equality, $\langle\dot{I}_{\{\rho\},Y}\rangle=d_{t}\mathcal{S}^{gb}(\mathcal{P}_{\{\rho\},Y})-d_{t}\mathcal{S}^{gb}(\mathcal{P}_{\{\rho\}})-d_{t}\mathcal{S}^{gb}(\mathcal{P}_{Y})$. The fluctuation theorem follows trivially by modifying eq. 54 that incorporates stochastic mutual information eq. 74, or equivalently replacing the reference state entropy with the reference conditional state entropy $S_{E}^{state}\to S_{E}^{state|Y}=-\log{(\mathcal{P}_{\{\rho\}|Y})}$. We have briefly highlighted the procedure to incorporate information thermodynamics for the non-reciprocal systems. The rest is rather trivial.

### v3

@@ Appendix A Microscopic excess and house-keeping decomposition of the EPR
The total mean microscopic bulk EPR can be further divided using another physical aspect of the system. For instance, a state variable is often easier to track compared to the underlying thermodynamic energy functional. One particular case arises when the temporal dynamics of the microstate lead to a steady-state probability distribution $P_{\{N\}}^{ss}$ for the microstate $\{N\}$. The mean bulk EPR ($\langle\dot{\sigma}^{\mathcal{R}}\rangle+\langle\dot{\sigma}^{\mathcal{D}}\rangle$) can be decomposed into two components, called excess and housekeeping EPR, using the steady state of the dynamics [107, 238, 239, 20, 104, 105, 237]. This decomposition is commonly referred to as the Hatano-Sasa (HS) decomposition of the total mean EPR. The mean excess HS EPR is denoted by $\langle\dot{\sigma}^{ex}_{hs}\rangle$, with the exact expression given by:
@@
Using the master equation eq. 7, the mean microscopic excess HS EPR satisfies $\langle\dot{\sigma}^{ex}_{hs}\rangle=\sum_{\{N\}}d_{t}P_{\{N\}}(t)\ln{\left(P^{ss}_{\{N\}}(t)\right)}$. The time dependence of the steady state indicates the validity of the definition in eq. 74 for non-autonomous dynamics implemented through thermodynamic work.
The exact expression for the microscopic housekeeping EPR reads:
@@
Positivity of excess and housekeeping EPR has been demonstrated [238, 239, 104, 105, 237]. By combining the rate of change of the Gibbs entropy term $d_{t}s^{gb}$ with the mean excess HS EPR $\langle\dot{\sigma}^{ex}_{hs}\rangle$, one defines the non-adiabatic mean EPR $\langle\dot{\sigma}^{na}\rangle$. Its closed-form expression, $\langle\dot{\sigma}^{na}\rangle=\langle\dot{\sigma}^{ex}_{hs}\rangle+d_{t}s^{gb}$, reads:
@@
The non-adiabatic mean EPR $\langle\dot{\sigma}^{na}\rangle$ quantifies microscopic EPR due to a mismatch between the instantaneous probability distribution $\{P_{\{N\}}(t)\}$ and the steady-state distribution $\{P^{ss}_{\{N\}}(t)\}$. In steady state, the instantaneous probability distribution satisfies $\{P_{\{N\}}(t)\}=\{P^{ss}_{\{N\}}(t)\}$, so that the non-adiabatic mean EPR $\langle\dot{\sigma}^{na}\rangle$ vanishes. Consequently, in steady state, the total mean microscopic EPR $\langle\dot{\sigma}\rangle$ is completely determined by the mean housekeeping EPR, $\langle\dot{\sigma}^{hk}_{hs}\rangle=\langle\dot{\sigma}\rangle$.
@@
Reorganizing the definitions of the mean microscopic excess HS EPR and the microscopic non-adiabatic EPR shows that they are related to the rate of change of the stochastic Massieu potential. In particular, $\langle\dot{\sigma}^{ex}_{hs}\rangle=d_{t}\langle\ln{(P^{ss}_{\{N\}}(t))}\rangle$ and $\langle\dot{\sigma}^{na}\rangle=-d_{t}\langle\ln{(P_{\{N\}}(t))/(P^{ss}_{\{N\}}(t))}\rangle$, where the time derivative is defined as $d_{t}\langle*\rangle=\sum_{\{N\}}(*)d_{t}P_{\{N\}}$. The same convention was previously used to define the second law of thermodynamics. Thus, $\ln{(P^{ss}_{\{N\}}(t))}$ and $\ln{(P_{\{N\}}(t))/(P^{ss}_{\{N\}}(t))}$ are stochastic Massieu potentials whose temporal variations are $\langle\dot{\sigma}^{ex}_{hs}\rangle$ and $\langle\dot{\sigma}^{na}\rangle$, respectively.
@@
The Kullback-Leibler divergence is defined as $D^{KL}_{ss}(P(t)||P^{ss}(t))=\sum_{\{N\}}P_{\{N\}}(t)\ln{(P_{\{N\}}(t)/P^{ss}_{\{N\}}(t))}$ [99, 100, 102, 103, 101, 104, 105, 106]. We introduce a shorthand notation $D^{KL}_{ss}(t)=D^{KL}(P(t)||P^{ss}(t))$, where the superscript indicates the reference distribution. Thus, $\langle\dot{\sigma}^{na}\rangle=-d_{t}D^{KL}_{ss}(P(t))$. Physically, this relates the non-adiabatic EPR to the relaxation of the Kullback-Leibler divergence between the instantaneous and steady-state distributions. Plugging this into eq. 75 gives $\langle\dot{\sigma}^{hk}_{hs}\rangle=\langle\dot{\sigma}\rangle|_{P^{ss}}$, showing that the housekeeping EPR equals the total microscopic EPR in steady state, while the microscopic adiabatic EPR vanishes. For a system without external non-conservative or non-reciprocal forces, $\langle\dot{\sigma}^{hk}_{hs}\rangle=0$ trivially holds. The microscopic EPR due to relaxation is fully quantified by the rate of change of the Kullback-Leibler divergence, $\langle\dot{\sigma}\rangle=-d_{t}D^{KL}_{ss}(t)$.
A stronger bound on the microscopic entropy production is obtained by exploiting the positivity of $\langle\dot{\sigma}^{na}\rangle$ and its Kullback-Leibler representation [106]. In particular, $\langle\sigma^{na}\rangle=D^{KL}_{ss}(0)-D^{KL}_{ss}(\tau)$, where $\tau$ is the relaxation time. Using the triangle inequality [106, 103], this can be further simplified as $\langle\sigma^{na}\rangle\geq D^{KL}(P(0)||P(\tau))$, which is a tighter bound than the second law of thermodynamics due to the non-negativity of the KL divergence.